OpenStudy (anonymous):

how do i use this formula to evaluate the sum of the even terms (equation inside)

OpenStudy (anonymous):

\[sum_{n=1}^{infty} 1/(2n)2\]

OpenStudy (anonymous):

well I have no idea what that is. It didn't work

OpenStudy (amistre64):

\[\sum_{n=1}^{\infty}\ \frac{1}{(2n)^2}\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{1}{(2n)^2}\]

OpenStudy (anonymous):

The script must be failing i can't read if i've laid that out correct

OpenStudy (amistre64):

write out a sequence and see if you can construct a new summer from it

OpenStudy (anonymous):

Now it's working.. satellite73 that is the question.. how do i go about solving it?

OpenStudy (anonymous):

\[1/4+1/16+1/36+1/64+1/100\]

OpenStudy (jamesj):

Actually, the way you deal with this in contemporary mathematics is with an integral in the complex plane. Easyish to explain with a blackboard where I can draw diagrams easily, annoying to do it here. For the record, the answer is pi^2/24.

OpenStudy (anonymous):

hrmm... Thanks james.. I really need to find a way to show working.

OpenStudy (amistre64):

\[\sum_{j=1}^{\infty}2k\ \sum_{n=k}^{\infty}\ \frac{1}{(2n)^2}\] hmmmmm

OpenStudy (jamesj):

For a classical proof, have a look at this ,around page 7 onwards

OpenStudy (jamesj):

http://www.maa.org/pubs/Calc_articles/ma045.pdf

OpenStudy (jamesj):

the result sum n = 1 to infty of 1/n^2 is called Euler's formula. That's what is shown in the note I just linked to. It gives pi^2/6 It should be straight-forward to imitate that derivation for your summand 1/(2n)^2

OpenStudy (jamesj):

Actually, it's even easier. Just pull the 1/2^2 = 1/4 outside. So what you just need to do is prove Euler's formula using what arrows you currently have in your analytical quiver.

OpenStudy (jamesj):

anyone there? Anyway ... time for brunch!